OpenStudy (anonymous):

Solve for x: ln(3x)^2=16.

OpenStudy (goformit100):

differentiate on both side w.r.t. x. .

OpenStudy (anonymous):

\[\ln(3x)^2=16\\ 2\ln 3x=16\\ \ln 3x=8\\ 3x=e^8\\ x=\frac{e^8}{3}\]

OpenStudy (anonymous):

Thanks!